OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 1) 0, -1, 1, 0, 2 2) 0, 1, 1, 2, 3 3) 0, 1, 2, 2, 3 4) 0, 1, 1, 2, 2

OpenStudy (perl):

$$\Large{ f(1)=0\\ f(2)=1\\ f(3) = f(2)+f(1) = 1 + 0\\ f(4) = f(3) + f(2) = ? } $$

OpenStudy (jdoe0001):

\(\large { \begin{array}{crllll} term&value \\\hline\\ 1&f(1)\to 0\\ 2&f(2)\to 1\\ {\color{brown}{ 3}}&f({\color{brown}{ n}}-1)+f({\color{brown}{ n}}-2)\\ & f({\color{brown}{ 3}}-1)+f({\color{brown}{ 3}}-2)\\ &f(2)+f(1)\\ {\color{brown}{ 4}}&f({\color{brown}{ n}}-1)+f({\color{brown}{ n}}-2)\\ & f({\color{brown}{ 4}}-1)+f({\color{brown}{ 4}}-2)\\ &f(3)+f(2)\\ \end{array} }\) notice perl 's line above, assuming you haven't fainted, seems like